How would this question be answered using formulas in Excel? Please show the formulas you use.A B C D E F G H I J K L M nm to 000 Question 2: In a random sample of 400 residents of Nassau County, 320 residents indicated
Answer #1

sample size sample proportion 320 400.0 0.8000 Ô x/n= std error =Se =v(p*(1-p)/n) = 0.0200 for 95 % CI value of z= margin of

a)

from above 95% confidence interval =0.761 , 0.839

b)

from above 99% confidence interval =0.748 , 0.852

(Note: for excel use highlighted formula:

for 95% confidence interval]

lower bound =(0.8-normsinv(0.975)*sqrt(0.8*(1-0.8)/400))

upper  bound =(0.8+normsinv(0.975)*sqrt(0.8*(1-0.8)/400))

for 99% confidence interval]

lower bound =(0.8-normsinv(0.995)*sqrt(0.8*(1-0.8)/400))

upper  bound =(0.8+normsinv(0.995)*sqrt(0.8*(1-0.8)/400))

)